OpenStudy (anonymous):

Find the solution to the equations. 2x - y = -3 x + y = 0 (0, 0) (-1, 1) (1, -1)

OpenStudy (anonymous):

just plug each ordered pair in until the equations are true, that is your answer

OpenStudy (anonymous):

(-1, 1)

OpenStudy (anonymous):

right! :)

OpenStudy (anonymous):

Thanks!:)

OpenStudy (anonymous):

no problem! message me for help anytime!

OpenStudy (anonymous):

2x-3=y x+y=x+2x-3=3x-3=0 3x=3 x=1 1+y=0 y=-1